### v1

# The Magnetic Ground State of Atacamite Cu2Cl(OH)3: The Crucial Role of Frustrated Zigzag Chains Revealed by Inelastic Neutron Scattering

Abstract
We report inelastic neutron scattering (INS) measurements on the magnetically frustrated sawtooth-chain compound atacamite Cu2Cl(OH)3 featuring inequivalent Cu(1) and Cu(2) sites. Transverse to the sawtooth chains, INS reveals two dispersive spin-wave modes and a gap of at least 0.75 meV. This behavior is rationalized within a zigzag-chain model of Cu(2) spins in an effective magnetic field of Cu(1) spins. The model is compatible with first-principles calculations and accounts for INS dispersions within linear spin-wave theory calculations. Our results reveal a unique case of an effective separation of energy scales between two differently oriented one-dimensional chains, with the zigzag-chain model being essential to fully characterize atacamite’s low-energy magnetism.
I I. Introduction
Frustrated spin systems offer the distinct potential to host novel exotic magnetic phases [1, 2, 3, 4] and have become a platform through which to explore fundamental magnetism. The exotic phases emerge from specific magnetic interaction models. For instance, the corner-sharing tetrahedra of pyrochlore lattices can produce spin-ice states [5, 6, 3], and the kagomé lattice is expected to host a quantum spin-liquid state [2]. Identifying experimental systems that realize these models allows for rigorous testing of theoretical predictions under real-world conditions.
Herbertsmithite (ZnCu3(OH)6Cl2) [7] is one of the leading candidates for an experimental realization of a quantum spin-liquid material with its geometrically perfect spin- kagomé lattice. It sits within the broader atacamite family of minerals [8], with the general formula (ZnxCu4-x(OH)6Cl2) [9] encompassing polymorphs with various ratios of Zn-Cu substitution.
Atacamite (Cu2Cl(OH)3) refers to the orthorhombic polymorph of this mineral family, with full Cu2+ occupation. The spin- Cu2+ sites class atacamite as a quantum-magnetic system, and experiments have characterized its geometric frustration effects [10]. Furthermore, atacamite exhibits a complex magnetic phase diagram, including an antiferromagnetic ground state at = 8.9 K [11] (with a magnetic propagation vector of = [, 0, ] [12] and with suppressed magnetic momeents of and for Cu(1) and Cu(2) sites, respectively.) and a sequence of metamagnetic transitions up to 30 T, including a plateau-like behvaior of the bulk magnetization above 31.5 T (H axis) [11]. However, the precise nature of these transitions and phases remained ambiguous.
Recently, a field-induced quantum critical point at 21.9(1) T (H c axis) was identified at the low-temperature endpoint of the long-range ordered antiferromagnetic phase [13]. Further, the results from Ref. [13] underpinned the scenario of a dimensional reduction to an effectively one-dimensional behavior occurring at the quantum critical point as the residual 3D exchange couplings are overcome by the magnetic field energy. In this work, we explore the low-energy spin-wave excitations in zero applied field, allowing us to probe in detail the residual exchange couplings in atacamite responsible for the long-range magnetic order below = 8.9 K and below the critical magnetic field.
Regarding the magnetic interactions driving the ground-state magnetism of atacmite, recent density functional theory (DFT) calculations indicate that the dominant exchange interactions form a weak 3D network of antiferromagnetic sawtooth chains [11]. The five largest exchange pathways are shown in Fig. 1, with the sawtooth chain running along the direction. Interest has followed atacamite not only for its compositional similarities to Herbertsmithite, but also due to the unmatched opportunity it offers to experimentally investigate the complexities of a low-dimensional quantum spin system within the intensely-studied sawtooth chain model [14, 15, 16, 17, 18, 19, 20, 21].
Although the sawtooth chain is the dominant coupling motif in atacamite, DFT calculations also reveal a weaker perpendicular chain formed by competing nearest and next-nearest interactions [11]. This is the S= zigzag spin chain [22, 23, 24, 25] (hereafter referred to as simply the ‘zigzag chain’ to avoid confusion with our exchange-label naming convention), with Hamiltonian
Here, and represent the nearest- and next-nearest-neighbor interactions, respectively, on a spin chain with an even number of spin sites and spin- operators at spin site .
The zigzag spin chain exhibits diverse magnetic phases dictated by the ratio of [26, 24] owing to the complexity of the competing interactions. These include gapped and gapless spin states, spiral orders, and frustration-driven quantum phases such as spin-nematic states [24, 25] and spin-stripe phases. These states have also been revealed for material realizations of the zigzag chain [27, 28, 29].
In this investigation, we demonstrate that atacamite’s zigzag chain plays a previously overlooked role in its magnetic behavior at low temperatures. We present the first investigation of atacamite’s spin-wave excitation spectrum using inelastic neutron scattering (INS) within its ordered phase to directly probe its magnetic interactions. We find that the zigzag chain is crucial for modeling atacamite, and, when combined with a mean-field treatment, an effective-model Hamiltonian for atacamite is produced that agrees with INS measurements in the low-energy regime. Furthermore, this agreement stands in contrast to the poor match obtained using the full DFT-derived model, whose dominant sawtooth chain is of a much higher energy scale than the zigzag chain. An effective separation of energy scales between the two chains is thus presented as a puzzling characteristic of a material realization of a frustrated quantum magnet.
The compatibility between a magnetic model and INS measurements can only be justified within a robust theoretical framework. Linear spin-wave theory (LSWT) stands out as a straightforward, common [30], and reliable method for making quantitative comparisons between a magnetic Hamiltonian and INS experiments in long-range ordered spin systems. The surprising success of LSWT in its application to quantum systems is rooted in the abundance of semiclassical behaviors [31], even in the limit [32]. The LSWT calculations within this work were performed using the Sunny package [33] where the magnetic structure refined from neutron diffraction data [11] was chosen for the spin structure input.
II II. Experimental Methods
LSWT and INS measurements are often used together to validate first-principles calculations and refine empirical magnetic models, including for frustrated (quantum) spin systems [5, 34, 35, 36, 37, 38]. Single-crystal neutron scattering measurements of atacamite’s spin-wave spectrum were performed at the Australian Nuclear Science and Technology Organisation (ANSTO) on the Pelican [39] time-of-flight INS spectrometer with the sample rotated in the H0L plane. Pelican was configured with 4.69 Å neutrons and an approximate resolution of 0.135 meV at the elastic line.
The single-crystal INS data represents a four-dimensional dataset, comprising three spatial dimensions and one energy dimension. Although aligned in the H0L plane, the 0K0 direction could be measured out-of-plane up to due the large detector height at Pelican of 1 m, resulting in r.l.u. of measurable range in K.
Data was collected for the sample at 1.5 K, 6 K, and 20 K in a closed-cycle helium cryostat, and was measured across a sample rotation of in intervals. A vanadium sample was used to normalize the detector efficiencies and an empty cryostat scan was used for background subtraction.
III III. Results
The INS data along the H-reciprocal-lattice direction, shown in Fig. 2(a), is a 2D cut where neutron counts have been integrated within the interval [–2.5,2.5] along L, and [–0.2,0.2] along K. The data within the volume along L is relatively non-dispersive, and the intensity is distributed over a broadened energy range, as shown in Section VI of the supplemental material [41]. The measurements probe the low-energy region of atacamite’s spin-wave excitations with an upper-energy threshold of 2.5 meV.
Along the H direction, two distinct, dispersive, energy-offset spin-wave modes are observed. A linear energy cut at H = 0.5 r.l.u. is shown in Fig. 3 and is taken with the same integration widths in L and K as the H-dependent spectra in Fig. 2(a), with an integration width of 0.04 r.l.u. in H. These data reveal two strong peaks in intensity – one at 1.7 meV and one at 1.3 meV. The lower-energy 1.3 meV peak is broad, and the neutron count doesn’t reach the background level at the magnetic zone center until approximately 0.75 meV, as shown in Fig. 4 (see Fig. S10 [41] for more details on the gap quantification). Atacamite thus has an energy gap of at least 0.75 meV along the H00 direction, but, due to instrumental broadening, could be as high as the peak center of meV, which would be compatible with the gap extracted from heat capacity measurements [13].
LSWT calculations performed using the five largest exchange interactions yielded by previous DFT results [11], which includes both the sawtooth and zigzag chains, fail to accurately reproduce all characteristics of these INS measurements. Specifically, spin-wave excitations appear at higher energies than are observed in the INS measurements, the asymmetry of the mode intensities do not match, and the lower-energy mode is not gapped (for further LSWT details for the full-DFT coupling network, see Figs. S5 and S6 [41]).
These shortcomings demonstrate that the full-DFT model does not adequately capture the magnetism governing atacamite’s low-energy excitations within a semiclassical approach. The discrepancies in the full-DFT model may stem from a the simplified treatment of quantum effects in LSWT, and could be alleviated in a fully quantum-mechanical treatment. Unfortunately, frustration and the three-dimensional nature of the full model make such treatment unfeasible. As an alternative, for a given energy scale, we can construct an effective model, which treats a subset of relevant exchanges exactly, while other exchanges enter the model in a form of an effective field [22, 23, 24, 25]. In atacamite, we achieve this by a mean-field treatment of the antiferromagnetic zigzag chain model along the axis using the Hamiltonian
Here, is the nearest-neighbor interaction along , is the next-nearest-neighbor interaction for the Cu(2) spins sites as defined in Fig. 1, and is an effective staggered mean-field term which mimics the coupling to Cu(1) spins that are part of the sawtooth chains. The 4-fold periodicity of the staggered field has been constructed from observing the same periodicity in the ordering of the Cu(2) spin chain itself. This model is represented in Fig. 5.
The parameters of this model have been refined through least-squares fitting of LSWT calculations to the INS data. The refinement arrives at = 0.90 meV, = 1.91 meV, and = 0.255 meV, giving a characteristic / ratio of 2.12. Classical LSWT calculations of the antiferromagnetic zigzag chain with ratios above 1 also match exact diagonalization calculations (see Fig. S11 in the supplemental information [41]), suggesting quantum effects are negligible for these calculations.
The refined LSWT scattering function calculation is shown in Fig. 2(b). The result shows a remarkable qualitative match with experimental data along the H direction, namely the spin-wave amplitudes, the relative intensities of the dispersion, asymmetry in the dispersion intensities, the weakening intensity towards the magnetic zone boundary, the spin-wave frequencies, and, crucially, the spin-wave gap. A 1D-cut of the calculated spin-wave spectra further confirms an excellent match with the experimental peak positions and gap size, as shown in Fig. 3.
The zigzag-chain model also provides critical insight into atacamite’s gapped ground state. Theoretical considerations of the unmodified zigzag chain [44, 22] suggest that a / ratio of 0.5 would be required to produce a gap within the measured range of our INS data. Atacamite’s much larger refined ratio of 2.12 would produce a gap in the order of only 0.1 meV [22]. This, and our observations that directly and independently controls the gap size, suggest that the gap is instead primarily a result of interchain interactions modeled by the mean-field term. Consequently, interchain frustration, particularly via connecting the transverse chains, must be the dominant mechanism driving the observed gap. This is consistent with two observations. One – that the spin configurations of the Cu(1) and Cu(2) sublattices are near-perpendicular, a sign that the zigzag chain has adopted a compromised non-colinear configuration with respect to to minimize frustration. And two – the Cu(2) spins have four-fold commensurate periodicity.
Without frustration, a zigzag chain is expected to adopt an incommensurate spiral order which is nearly of four-fold periodic [22]. The fact that the ground state is exactly four-fold commensurate implies that weak quantum fluctuations have selected this spin state through order-by-disorder from a frustration-induced classically degenerate set. This breaks symmetry, explaining the large observed gap [45].
While the LSWT-refined parameters provide a compelling and accurate model of the spin-wave excitations, they differ significantly from the initial DFT-calculated exchange values. The nominal DFT values were calculated using the generalized gradient approximation (GGA) method [46] with a Coulomb repulsion -value of 8.24 eV [11]. These are = 0.11 meV and = 1.35 meV, with a / ratio of 12.27, which fail to provide agreement with INS measurements [41].
However, the value has a drastic influence on the value and even the sign of [11], and hence also the / ratio, which plays the central role for the ground state. To inspect this difference closely, we performed new DFT calculations using the full-potential code FPLO version 22 [47]. We used the GGA supplemented with the Coulomb repulsion which we varied in the range between 7 and 9 eV, and the Hund exchange of 1 eV. Further details of the calculations are provided in Section IV the supplemental information [41]. The GGA+ energies were mapped onto a classical Heisenberg model. In this way, we found that the LSWT-determined ratio is reproduced by the interpolated values of = 1.40 meV and = 0.66 meV at eV. The exchange-coupling values for initial DFT calculations from the literature, our new DFT calculations, and the fitting results for the zigzag chain are shown in Table 1.
| Exchange | Previous DFT [11] | New DFT | Zigzag-Model LSWT Refinement |
| label | (meV) | (meV) | (meV) |
| 0.11 | 0.66 | 0.98 | |
| -0.83 | -0.83 | – | |
| 8.79 | 9.39 | – | |
| 28.97 | – | – | |
| 1.35 | 1.40 | 1.94 | |
| – | – | 0.26 |
IV IV. Discussion
Our results have revealed the critical nature of in modeling atacamite, and INS has allowed us to refine its highly -sensitive value. These results, however, do not directly cast doubt on the validity of the previous DFT approach; the previous calculations correspond to the energy scale of the Curie-Weiss temperature [11], their dominant couplings agree with the Goodenough-Kanamori-Anderson rule [48, 49, 50] by association with atacamite’s largest Cu-O-Cu superexchange pathways, and they contain the very zigzag chain geometry used for our model.
Instead we make three key observations: Firstly, the zigzag chain plays an important role in atacamite’s magnetism not previously explored. The necessity of the zigzag chain to reproduce the ordered-phase spin-wave dispersion at 1.5 K indicates that the zigzag-chain couplings are essential to the magnetic ordering behavior in atacamite. Atacamite’s measured ground state behaves as a zigzag chain in a weak staggered field. It is therefore not enough to simply describe atacamite as a one-dimensional sawtooth chain system when discussing its long-range ordered phase.
Secondly, the energy scales between the zigzag and sawtooth chain are effectively separated, reflecting different roles in atacamite’s spin dynamics. Although all isotropic couplings can be deduced simultaneously through DFT, low-energy spin-wave calculations performed with the full coupling scheme are unable to fully reproduce INS observations. Meanwhile, at temperatures far above TN = 8.9 K, the Curie-Weiss temperature derived from inverse susceptibility aligns with the dominant energy scale of the sawtooth chain. Additionally, magnetic order is suppressed near the energy scale of the zigzag chain [13], suggesting it governs atacamite’s ordering dynamics.
This effective energy scale separation may be explained by frustration effects in atacamite. The refined mean-field term, meV, is significantly smaller than the dominant interchain coupling predicted by DFT ( 9 meV), indicating that frustration suppresses the effective interchain coupling. Driven by , this frustration not only inhibits long-range order by competing within the dominant sawtooth chain but also manifests itself through the order-by-disorder effects discussed earlier. As a result, atacamite’s magnetic ground state is more sensitive to fluctuations, which both enhances the importance of subtle Hamiltonian terms, such as anisotropic interactions, and effectively separates the system into two distinct 1D coupling schemes. These same features also make it particularly challenging to construct a complete and precise model based on direct-exchange DFT parameters alone. Small inaccuracies in subtle interactions can become amplified in a fluctuation-influenced system, where semiclassical calculations may respond sensitively to such details.
Lastly, is extremely sensitive to , which means a highly-precise determination of is crucial for a model of atacamite. Through LSWT refinement against INS measurements, we have provided a significant and thorough refinement of the value required to accurately describe atacamite in the low-temperature regime associated with long-range order.
V V. Conclusions
LSWT calculations and INS measurements have revealed key insights into atacamite’s magnetism in its long-range ordered phase. The mean-field-modified zigzag-chain model accurately reproduces the periodicity, intensity asymmetry, dispersion, and spin-wave gap observed in the INS data. The effective mean-field term estimates the frustrated interchain interaction strength and provides insight into the spin-wave gap through an order-by-disorder mechanism.
In contrast, while prior DFT calculations for the dominant direct exchange couplings are consistent with high-temperature experimental results, they fail to account for the low-energy spin dynamics below the ordering temperature in a semiclassical LSWT treatment. This suggests that the full isotropic-exchange DFT model may be missing the subtleties required to capture the low-energy magnetic dynamics in a frustrated magnet. Frustration also enhances quantum effects, suppressing weaker interactions such as interchain couplings. This has led to an effective energy scale separation between two distinct 1D-chain models. Nevertheless, by employing an effective model of a field-moderated zigzag chain, we successfully capture the essential features of the experimental dispersion relations and provide a reliable description of atacamite’s low-energy magnetism. This way, we have uncovered the frustrated interchain exchange in atacamite leading to long-range magnetic order below its Néel temperature, and where the magnetic state is selected through order-by-disorder. In the light of the recent experimental results in high magnetic fields [13], it seems to be this frustrated interchain coupling network which is effectively broken up when the Cu(2) spins are fully field polarized in magnetic fields beyond the quantum critical point. Numerical results here supported the scenario of an (effective) dimensional reduction of the spin system upon which long-range magnetic order is suppressed.
Through establishing the exchange interactions outside the sawtooth chain in more detail, an avenue is now opened for building a clearer picture of the field-induced transitions which depend on the spin dynamics connected to these subtler interactions.
VI Acknowledgments
This research used the Pelican time-of-flight spectrometer for proposal number P8226 at ACNS, a part of ANSTO, the Australian Nuclear Science and Technology Organisation. J. L. Allen would like to thank AINSE Limited for providing financial assistance (Award - PGRA) to enable this research. The authors acknowledge and thank A. U. B. Wolter from The Leibniz Institute for Solid State and Materials Research for her extensive contributions to atacamite research, and for constructive discussions while preparing the manuscript.